1/18c - 1/36 + 1/6c = 1/6 + c

Answer by rfer(16322)

1/18c-1/36+3/18c=5/36+18/18c
-14/18c=6/36
-7/8c=1/6
c=1/6*-8/7
c=-8/42
c=-4/21